OpenStudy (jack1):

no rush, just a brain teaser to keep u all sharp (psh, like u need it brainaics!): If you were at a party, and everyone shook hands with every other person, and there were 66 handshakes in total... then how many people were at the party? I'm off to a concert, will post the answer tomoz, slaters all and good luck ;)

OpenStudy (anonymous):

There is n+1 people and the number of handshakes is 1+2+3+4+5+6... The sum is therefore \[\frac{ n(n+1) }{ 2 }\] Each handshake is counted twice so we divide by 2 We solve that = 66 and get n(n+1)=132 n^2+n-132=0 and solve for the quadratic equation (n+12)(n-11)=0 Since the number of people cannot be -12 then it is 11 So the final answer is 11 I think

OpenStudy (***[isuru]***):

nope.. u r so close... but nt close enough :-) if there were "n" amount of people and u were one of them u have to do (n - 1) handshakes then the other person will have to do ( n-2) ... and so on until the amount of handshakes will be 1 so... this is an arithmetic progression.. 1 , 2, 3, 4, ...... ( n-1) and (n-1) is the "n" th term... so... Sn = n( a + l) /2 66 = (n-1)( 1 + n-1 ) /2 132 = n(n-1) n^2-n-132=0 (n-12)(n+11)=0 so... the amount of people is 12 not 11 .....